# Overestimation, Overfitting, and Plasticity in Actor-Critic: the Bitter Lesson of Reinforcement Learning

### v1

@@ 4.4 Correlation of Overestimation, Overfitting and Plasticity metrics with Performance
This study analyses the relationships between Overestimation, Overfitting, Plasticity, and model performance. Overestimation is quantified as an approximation error, overfitting as the ratio of TD error on the validation set to TD error on the training set. Expressing Plasticity loss is challenging, so we utilize proxy metrics, including the percentage of Dormant neurons, representations rank, gradient norm, and parameters norm.
We employ a Spearman correlation matrix to scrutinize these dependencies. This statistic is chosen because we observe non-linear yet monotonic dependencies between the mentioned metrics. We employ it on the data from all performed experiments, i.e., form runs with different combinations of interventions. Moreover, we do not have a division into RR=2 and RR=16, only the results from both setups are combined, and analyses are made on them. Overestimation, gradient norm, and parameters norm are analyzed in a logarithmic scale for precision. We exclude metric pairs where the p-value of correlation is above $5\%$ by whitening tiles in the correlation matrix.
Figure 5: Spearman correlation matrix for explanatory metrics on DMC benchmark. Blank spaces are correlations that do not meet the p-value.
Figure 6: Spearman correlation matrix for explanatory metrics on MetaWorld benchmark. Blank spaces are correlations that do not meet the p-value.
@@
Figures 5 and 6 depict Spearman correlation matrices for DMC and MW benchmarks, respectively. Notably, overestimation exhibits the strongest correlation with agent returns on both benchmarks, offering insights into the findings of previous sections regarding the limited robustness of critic regularization methods designed to minimize overestimation.
Interestingly, as shown in Figure 7, layer norm and spectral norm effectively mitigate approximation errors, outperforming CR methods. Further investigation on the DMC benchmark (Figure 5) uncovers a strong negative correlation between the percentage of dormant neurons and performance, closely tied to the rank of representations on the critic’s penultimate layer.
Moreover, on the MW benchmark (Figure 6), the second-best correlated metrics with performance are the critic gradient norm and the critic parameters norm. This aligns with the results from section 4.2, highlighting the significant performance boost provided by spectral norm and weight decay, particularly in the RR=16 setup. An in-depth analysis of how RR increases the negative correlation of gradient norm and return can be found in sections D.2 and D.3 of Appendix.
@@
Spearman correlation in dog environments:
To deepen analysis from the previous paragraph, we investigate the relations between variables separately for every environment. Interestingly, for the DMC benchmark, we find that overestimation is not a good predictor for dog environments where we observe the highest values of overestimation (Figure 8). We hypothesize that overestimation becomes a good predictor of performance only when more fundamental issues, such as plasticity, are mitigated, indicating a multifaceted learning problem in harder environments.
As an observation that supports this hypothesis, we refer to the critic gradient norm, which exhibits the most monotonic relation with the return in dog environments, as indicated by Spearman correlation (Figure 8). The high gradient norm directly points to high curvature of the loss landscape, which, as indicated by (Lee et al., 2023), describes low input plasticity. For this reason, layer norm primarily smoothens the activation distribution and plays a critical role in making SAC work in dog environments. Enviroments from the MW benchmark also encounter challenges with high curvature loss landscapes (as indicated by the Spearman correlation between gradient norm and return). This suggests a resemblance between MW and DMC dog environments.
@@
Environments can differ significantly in terms of explanatory metrics correlations, even those from the same benchmark suite. Techniques like layer or spectral norm and resets are particularly effective in mitigating overestimation also compared to methods specifically designed for that purpose. The negative correlation of the critic’s gradient norm and return becomes more apparent in challenging environments and mainly in a high replay ratio regime.
@@ 5 Related Works
Regularization schemes have proven effective in enhancing deep reinforcement learning methods. Neural network regularizations, such as Spectral Norm (Gogianu et al., 2021; Bjorck et al., 2021), Layer Norm (Bjorck et al., 2021; Ball et al., 2023), or weight decay (Liu et al., 2020), have yielded significant improvements in results. In contrast, periodic resets of critic weights were proposed by (Nikishin et al., 2022). An ensemble approach was also suggested to address the challenge of determining the optimal number of gradients per environment step, where decisions are based on the validation TD error (Li et al., 2022).
@@ 7 Conclusions
This study explored different common RL design choices, as well as interactions thereof, evaluating their impact on agent learning. Specifically, we consider three types of regularization families: critic regularization (motivated by value overestimation); network regularization (motivated by model overfitting); and plasticity regularization (motivated by plasticity loss). Our analysis revealed that generic network regularization methods such as layer normalization, especially when paired with full-parameter resets, can have vastly greater impact on the final performance than domain-specific RL approaches. To this end, the same network regularization methods can lead to strong performing policies on domains previously solved by model-based agents, such as the dog domain. Furthermore, we studied a variety of metrics which were shown to co-occur with deterioration of learning in low and high replay regimes. Our analysis revealed the complex interactions of the considered metrics with agent performance. Surprisingly, we found that interventions motivated by a specific problem, for example overfitting, can have a pronounced impact on the metrics associated with overestimation or plasticity. Finally, we found that the effectiveness of considered critic, network and plasticity regularization techniques is not only highly dependent on the simulation benchmark, but also type of simulated task. The most prominent example is Clipped Double Q-learning, a technique used in majority of modern actor-critic algorithms, is effective in DMC locomotion tasks, but leads to significant performance deterioration on the MetaWorld manipulation tasks. To this end, we highlighted the need of testing new algorithms on a diverse set of tasks, preferable stemming from more than one suite.
@@
Appendix A Derivations - Variance
Appendix B Details of experiments
Results reporting Interquartile mean (IQM) are based on 500 boostrapping points as calculated by rliable package (Agarwal et al., 2021). The final performance is defined as average of last 10 policy evaluations.
Appendix C Architecture details
@@
In the scenario involving Layer Norm, it is applied to each hidden layer (Li et al., 2022; Ball et al., 2023), while spectral norm is applied exclusively to the last hidden layer (Gogianu et al., 2021; Li et al., 2022). Weight decay is uniformly applied across all layers (Li et al., 2022). It’s important to note that all network regularizations are exclusively applied to the Critic network.
C.1 Hyperparameters
All hyperparameters are taken from original papers introducing given intervention.
@@ C.1 Hyperparameters
Appendix D Further Experiments
D.1 Third-order marginalization
@@
In a scenario of high replay ratio, the resets are the most important interevention. They work best in combination with Layer norm, but other neural network regularization methods combined with resets are at the forefront regarding performance in the rr=16 scenario (see top of the Figure 9). We present more results in Figures 20 and 21.
D.2 Gradient Norm Analysis
@@
A similar phenomenon can be observed on the DMC benchmark, but layer norm proves more robust in mitigating exploding gradients than Spectral Norm. Interestingly, training on very complex environments such as Dog causes enormous gradient explosion, even in a small replay ratio regime 11.
Figure 10: IQM gradient norm of first-order results. The IQM is calculated based on the average of the last ten evaluation points in each run, not the last evaluation point. Colors indicate hierarchy on the plot, not specific names.
Figure 11: IQM gradient norm of first-order results. The IQM is calculated based on the average of the last ten evaluation points in each run, not the last evaluation point. Colors indicate hierarchy on the plot, not specific names.
D.3 Regression and Spearman correlation analysis
@@
In figures 12 and 13, we observe consistent negative coefficients, except for the Coffee environment for overestimation and return. Interestingly, for the MetaWorld benchmark, most environments exhibit a symmetrical degree of overestimation and underestimation as two distinct clusters of well-performing runs are mirrored by the Y axis. For the DMC environments, we observe a much higher amount of runs in which overestimation is present.
For the gradient norm, we observe a higher value of this metric for runs with RR=16 both for DMC and MW benchmarks (Figures 14 and 15. There is consistency in coefficients for MW; however, in DMC, we observe that only four out of 7 environments exhibit negative signs of the coefficient. However, for most of the extremely high values of gradient norm, we systematically observe low returns regardless of the environment.
Dormant neurons, presented in Figures 16 and 17, clearly correlate negatively with the return, especially for DMC environments. However, there are cases, such as the Reach environment from MetaWorld, where a high percentage of dormant neurons benefit the agent, probably because this particular environment is especially easy.
Overfitting is the metric that exhibits the highest variance in coefficient signs across environments and benchmarks. As shown in Figures 18 and 19, the best-performing runs are located for low absolute values of overfitting. In addition, RR=16 clearly results in higher overfitting values.
@@
D.4 Best combinations of intervention performance plots
@@ D.5 Other
Figure 23: RR changes on metrics - MOVE TO APPENDIX

### v2

@@ 4.4 Correlation of Overestimation, Overfitting and Plasticity metrics with Performance
This study analyses the relationships between Overestimation, Overfitting, Plasticity, and model performance. Overestimation is quantified as an approximation error, overfitting as the ratio of TD error on the validation set to TD error on the training set. Expressing Plasticity loss is challenging, so we utilize proxy metrics, including the percentage of Dormant neurons (Sokar et al., 2023), representations rank (Kumar et al., 2020), gradient norm (Nikishin et al., 2022; Lyle et al., 2023), and parameters norm (Nikishin et al., 2022; Lyle et al., 2023).
We employ a Spearman correlation matrix to scrutinize these dependencies. This statistic is chosen because we observe non-linear yet monotonic dependencies between the mentioned metrics. We employ it on the data from all performed experiments, i.e., form runs with different combinations of interventions. Moreover, we do not have a division into RR=2 and RR=16, only the results from both setups are combined, and analyses are made on them. Overestimation, and gradient norm, and parameters norm are analyzed in a logarithmic scale for precision. We exclude metric pairs where the p-value of correlation is above $5\%$ by whitening tiles in the correlation matrix.
@@
In Figure 5, we investigate the relations between plasticity loss, overestimation and overfitting metrics and agents return, separately for every benchmark with special separation for Dog environments. Notably, overestimation exhibits the strongest correlation with agent returns on both benchmarks, offering insights into the findings of previous sections regarding the limited robustness of critic regularization methods designed to minimize overestimation.
Interestingly, as shown in Figure 6, layer norm and spectral norm effectively mitigate approximation errors, outperforming CR methods. However, further investigation on the DMC benchmark (Figure 5) uncovers a strong negative correlation between the percentage of dormant neurons and performance, closely tied to the rank of representations on the critic’s penultimate layer. What is more, the overestimation is not the best predictor for Dog environments where we observe the highest values of overestimation (Figure 5). We hypothesize that overestimation becomes a good predictor of performance only when more fundamental issues, such as plasticity, are mitigated, indicating a multifaceted learning problem in harder environments.
As an observation that supports this hypothesis, we refer to the critic gradient norm, which exhibits the most monotonic relation with the return in dog environments, as indicated by Spearman correlation (Figure 5). Analysing Figure 10 one can see, that Dog environments especially with high replay ratio experience exploding gradients. The high gradient norm directly points to high curvature of the loss landscape, which, as indicated by (Lee et al., 2023), describes low input plasticity. For this reason, layer norm primarily smoothens the activation distribution and plays a critical role in making SAC work in dog environments. Environments from the MW benchmark also encounter challenges with high curvature loss landscapes (as indicated by the Spearman correlation between gradient norm and return). This suggests a resemblance between MW and DMC dog environments.
Moreover, on the MW benchmark (Figure 5), the second-best correlated metrics with performance are the critic gradient norm and the critic parameters norm. This aligns with the results from section 4.2, highlighting the significant performance boost provided by spectral norm and weight decay, particularly in the RR=16 setup. An in-depth analysis of how RR increases the negative correlation of gradient norm and return can be found in sections C.2 and C.5 of Appendix.
Dog environments through gradients prism
@@
Figure 7: IQM gradient norm of first-order results. The IQM is calculated based on the average of the last ten evaluation points in each run, not the last evaluation point.
@@
There are distinctive correlations between plasticity loss, overestimation, overfitting metrics, and agent returns in various benchmark suites. These underscore the importance of considering environment-specific factors when assessing model performance and designing effective regularization strategies.
Techniques like layer or spectral norm and resets are particularly effective in mitigating overestimation also compared to methods specifically designed for that purpose. The negative correlation of the critic’s gradient norm and return becomes more apparent in challenging environments and mainly in a high replay ratio regime.
@@ 5 Related Works
Regularization schemes have proven effective in enhancing deep reinforcement learning methods. Neural network regularizations, such as Spectral Norm (Gogianu et al., 2021; Bjorck et al., 2021), Layer Norm (Bjorck et al., 2021; Ball et al., 2023), or weight decay (Liu et al., 2020), have yielded significant improvements in results. Notably, periodic resets of critic weights proposed by (Nikishin et al., 2022) constitute a strong baseline in robotics control. Several regularization schemes have been proposed to address the issue of discarding knowledge caused by fully resetting the critic network. Of particular interest is Shrink and Perturb (Ash & Adams, 2020) and L2 Init (Kumar et al., 2023). In both of these methods the benefit comes from the regularization towards the distribution of ”freshly” initialized weights. A concurrent work finds that using unit-ball normalization allows for learning with a high-replay ratio without full-parameter resets (Hussing et al., 2024). An ensemble approach was also suggested to address the challenge of determining the optimal number of gradients per environment step, where decisions are based on the validation TD error (Li et al., 2022).
@@ 7 Conclusions
This study explored different common RL design choices, as well as interactions thereof, evaluating their impact on agent learning. Specifically, we consider three types of regularization families: critic regularization (motivated by value overestimation); network regularization (motivated by model overfitting); and plasticity regularization (motivated by plasticity loss). Our analysis revealed that generic network regularization methods such as layer normalization, especially when paired with full-parameter resets, can have a vastly greater impact on the final performance than domain-specific RL approaches. To this end, the same network regularization methods can lead to strong performing policies on domains previously solved by model-based agents, such as the dog domain. Furthermore, we studied a variety of metrics that were shown to co-occur with the deterioration of learning in low and high replay regimes. Our analysis revealed the complex interactions of the considered metrics with agent performance. Surprisingly, we found that interventions motivated by a specific problem, for example, overfitting, can have a pronounced impact on the metrics associated with overestimation or plasticity. Finally, we found that the effectiveness of considered critic, network, and plasticity regularization techniques is not only highly dependent on the simulation benchmark but also type of simulated task. The most prominent example is Clipped Double Q-learning, a technique used in a majority of modern actor-critic algorithms, which is effective in DMC locomotion tasks, but leads to significant performance deterioration on the MetaWorld manipulation tasks. To this end, we highlighted the need to test new algorithms on a diverse set of tasks, preferably stemming from more than one suite.
Acknowledgements
Marek Cygan was partially supported by an NCBiR grant POIR.01.01.01-00-0433/20. Mateusz Ostaszewski was funded by the National Science Center Poland under the grant agreement 2020/39/B/ST6/01511. Michał Bortkiewicz was funded by the Warsaw University of Technology within the Excellence Initiative: Research University (IDUB) programme. This research was also supported by National Science Centre, Poland grant no 2020/39/B/ST6/01511 and grant no 2022/45/B/ST6/02817. This work was partially funded by the European Union under the Horizon Europe grant OMINO (grant number 101086321) and Horizon Europe Program (HORIZON-CL4-2022-HUMAN-02) under the project ”ELIAS: European Lighthouse of AI for Sustainability”, GA no. 101120237. We also gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2023/016783.
@@ Acknowledgements
Appendix A Details of experiments
Results reporting Interquartile mean (IQM) are based on 500 bootstrapping points as calculated by rliable package (Agarwal et al., 2021). The final performance is defined as the average of the last 10 policy evaluations.
Appendix B Architecture details
@@
In the scenario involving Layer Norm, it is applied to each hidden layer (Li et al., 2022; Ball et al., 2023), while the spectral norm is applied exclusively to the last hidden layer (Gogianu et al., 2021; Li et al., 2022). Weight decay is uniformly applied across all layers (Li et al., 2022). It’s important to note that all network regularizations are exclusively applied to the Critic network.
B.1 Hyperparameters
All hyperparameters are taken from original papers introducing the given intervention.
@@ B.1 Hyperparameters
Appendix C Further Experiments
C.1 Third-order marginalization
@@
In a scenario with a high replay ratio, the resets are the most important intervention. They work best in combination with Layer norm, but other neural network regularization methods combined with resets are at the forefront regarding performance in the rr=16 scenario (see the top of Figure 8). We present more results in Figures 22 and 23.
C.2 Gradient Norm Analysis
@@
A similar phenomenon can be observed on the DMC benchmark, but the layer norm proves more robust in mitigating exploding gradients than the Spectral Norm. Interestingly, training on very complex environments such as Dog causes enormous gradient explosion, even in a small replay ratio regime 10.
Figure 9: IQM gradient norm of first-order results. The IQM is calculated based on the average of the last ten evaluation points in each run, not the last evaluation point.
Figure 10: IQM gradient norm of first-order results. The IQM is calculated based on the average of the last ten evaluation points in each run, not the last evaluation point.
C.3 Comparison of ReDO to other plasticity-inducing methods
The ReDO method, as proposed by Sokar et al. (Sokar et al., 2023), is a technique for inducing network plasticity. It shares similarities with the full reset approach but involves more targeted interventions within the network. In particular, ReDo does not reset the full network; it only resets weights connected to dormant neurons. Specifically, incoming weights to dormant neurons are initialized as in full reset, but outgoing weights from dormant neurons are zeroed out, resulting in less severe network output changes.
Figure 11 presents results from Figure2 updated with the ReDo method for both MetaWorld and DMC environments for a high replay ratio regime. ReDo does not perform as well as resets in the Metaworld and DMC suites. However, both methods effectively reduce the critic gradient norm, overestimation and the number of dormant neurons for both Metaworld and DMC without dog benchmarks, as shown in Figure 12. In dog environments, we observed that ReDo was unstable for runs without SN or LN, and some runs crashed. We report results for the last ten timesteps before the crash for these runs. Interestingly, all methods except SN and LN cannot reduce the critic gradient norm, as shown in the bottom right plot in Figure 12.
Figure 11: Second-order results marginalizing critic regularization methods with ReDO.
Figure 12: IQM of performance, overestimation, dormant neurons, and gradient norm, of first-order results. The IQM is calculated based on the average of the last ten evaluation points in each run, not the last evaluation point.
C.4 Closer look on CDQL performance on Hopper and Quadruped environments
Figure 13 illustrates the outcomes of applying Clipped Double Q-learning (CDQL) (Haarnoja et al., 2018; Fujimoto et al., 2018; Hansen et al., 2022), a widely used and straightforward critic regularization method, across various environments. In the case of the hopper hop environment or some MW environments (bottom row), CDQL exhibits a detrimental effect on performance, with additional regularization techniques such as resets or layer normalization failing to alleviate this effect. Similarly, in the quadruped run task, CDQL demonstrates a comparable negative impact, although the application of additional regularization methods successfully mitigates it.
Figure 13: Examining the influence of CDQL on performance in the Hopper Hop and Quadruped Run environments within the DMC benchmark (top row), as well as the Push, Coffee, and Hand environments within the MW benchmark (bottom row).
C.5 Regression and Spearman correlation analysis
@@
In figures 14 and 15, we observe consistent negative coefficients, except for the Coffee environment for overestimation and return. Interestingly, for the MetaWorld benchmark, most environments exhibit a symmetrical degree of overestimation and underestimation as two distinct clusters of well-performing runs are mirrored by the Y axis. For the DMC environments, we observe a much higher amount of runs in which overestimation is present.
For the gradient norm, we observe a higher value of this metric for runs with RR=16 both for DMC and MW benchmarks (Figures 16 and 17. There is consistency in coefficients for MW; however, in DMC, we observe that only four out of 7 environments exhibit negative signs of the coefficient. However, for most of the extremely high values of gradient norm, we systematically observe low returns regardless of the environment.
Dormant neurons, presented in Figures 18 and 19, clearly correlate negatively with the return, especially for DMC environments. However, there are cases, such as the Reach environment from MetaWorld, where a high percentage of dormant neurons benefit the agent, probably because this particular environment is especially easy.
Overfitting is the metric that exhibits the highest variance in coefficient signs across environments and benchmarks. As shown in Figures 20 and 21, the best-performing runs are located for low absolute values of overfitting. In addition, RR=16 clearly results in higher overfitting values.
@@
C.6 Image-based DeepMind Control
We test whether the results achieved on proprioceptive state representation transfer to image-based control. To this end, we run 4 versions of the DrQ agent (Yarats et al., 2020):
  1. 1.
    Vanilla DrQ (DrQ)
  2. 2.
    DrQ with layer normalization on the critic network (DrQ + LN)
  3. 3.
    DrQ with full-parameter resets every 200k environment steps (DrQ + Res)
  4. 4.
    DrQ with both normalization on the critic network and full-parameter resets every 200k environment steps (DrQ + LN + Res)
We run these variations on 6 tasks from the DeepMind Control benchmark: Acrobot Swingup, Cheetah Run, Hopper Hop, Humanoid Run, Humanoid Stand, and Humanoid Walk. We run the humanoid tasks for 3mln frames and the other tasks for 1mln frames with a replay ratio of 1. We calculate the relationship between frames and environment steps according to the methodology presented in Yarats et al. (2020). We present the results in the table below.
Table 2: Final performance in image-based environments. 3 seeds per task.
DrQ DrQ + LN DrQ + Res DrQ + LN + Res
Acrobot Swingup $172.9\pm 22.1$ $87.6\pm 19.7$ $52.3\pm 11.2$ $88.8\pm 8.2$
Cheetah Run $727.4\pm 7.3$ $680.0\pm 2.3$ $715.6\pm 7.1$ $653.9\pm 2.1$
Hopper Hop $74.6\pm 34.9$ $116.3\pm 27.2$ $135.7\pm 24.1$ $167.3\pm 14.3$
Humanoid Stand $7.8\pm 0.2$ $8.1\pm 0.2$ $7.5\pm 0.4$ $7.8\pm 0.4$
Unfortunately, the humanoid agents were mostly unable to achieve non-random policies in the budget of 3mln frames. Interestingly, the proprioceptive results do not seem to directly transfer to image-based agents with a low-replay ratio. As such, we believe the image-based benchmark requires further studies.
C.7 Best combinations of intervention performance plots
@@ C.8 Other
Figure 25: Spearman correlation matrix for explanatory metrics on DMC benchmark (left) and on MetaWorld benchmark (roght plot). Blank spaces are correlations that do not meet the p-value.
Figure 26: Spearman correlation for different replay ratios.